Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-12-29

Observers: Charlotte Angus, Sadie Jones
TAT: Janet Chen, Lluís Galbany, Kate Maguire
Support team: Southampton

**  Problems with NTT electronics were reported to the TO prior to afternoon setup, telescope was inoperable until 00:13 UT  **

Calibrations:  unable to run day time calibrations, biases, spec arcs  (morn)

Conditions: Cloud throughout the night, wind stable at 5m/s. Disgusting seeing around 1.7" at start of night but improving to <1" by the end. 


UT               Target                        Type                          Comment
==============================================================================================
00:13         Issues with NTT electronics resolved. TO running telescope start-up procedures. 
00:55          AT2017bzc            followup                        Gr11, slit#1.5  Seeing ~1.7" at beginning of observation
01:48          AT2017bzc            followup                        Gr16, slit#1.5  Seeing ~1.5"
02:30          AT2017jei            classification                 SN-IIn   -13d
03:10          AT2017jan       followup                            Gr11 Seeing ~1.4"
04:05          AT2017jan       followup                            Gr16 Seeing ~1.4"
05:05          GD71               standard                            seeing 1"
05:35        SN2017iuk        followup                            Gr11 seeing 1"
06:34        SN2017iuk        followup                            Gr16 seeing <1"
07:17       AT2017iue        followup                            Gr13 seeing 1"
08:12       AT2017izu       classification                      Ia -1d      
08:20       AT2017jfw      classification                       Ia >+20d
08:35       L745a           standard                                 seeing 1"           
Comments